OpenStudy (anonymous):

can someone help me with this problem. Solve the equation. 6/(x+1)-1/3=1/2x+2

OpenStudy (anonymous):

Sure. Does your equation look like this? \(\dfrac{6}{x+1} - \dfrac{1}{3} = \dfrac{1}{2}x + 2\)

OpenStudy (anonymous):

\[(6divx+1)-(1\div3)=1\div(2x+2)\] \[(18-x-1)\div(3x+3)=1\div(2x+2)\] (17-x)(2x+2)=3x+3 \[solve and u get....29x-2^{2}-31=0\] finally....2x^{2}-29x+31=0